OpenStudy (masterchief):

Can someone create an equation with this information?: x: 5 y: 4 10 40 15 76 20 112

OpenStudy (bahrom7893):

let's see.. x1=5, y1=4 x2=10, y2=40 M = (40-4)/(10-5) = 36/5

OpenStudy (bahrom7893):

y-4=(36/5)(x-5) y = 36x/5 - 32

OpenStudy (bahrom7893):

now let's check if it works

OpenStudy (bahrom7893):

36*10/5 - 32 = 40 36*15/5 - 32 = 76.. etc, it works!

OpenStudy (masterchief):

y=36(10)/5 - 32 y=360/5 - 32 y=72-42 y=40 Thank you sir =)

OpenStudy (masterchief):

I didn't reply because my internet is messed up, it switches off and on, thx